Minnie and Helen are playing in the same hockey match. The probability of Minnie scoring a goal is 0.3. The probability of Helen scoring a goal is 0.4. What is the probability of both Minnie and Helen scoring a goal.

The probability of two events happening is equal to the probability of event 1 times the probability of event 2. Therefore, the probability of both Minnie and Helen scoring a goal is equal to the probability of Minnie scoring a goal times the probability of Helen scoring a goal.
0.3 x 0.4 = 0.12

MA

Related Maths GCSE answers

All answers ▸

Solve the simultaneous equations 3x + 2y = 4 and 4x + 5y = 17


There are 10 boys and 20 girls in a class. The mean mark for the class is 60, the mean mark for the girls is 54. Work out the mean mark for the boys.


express (2x^2-3x-5)/(x^2+6x+5) in the form (ax+b)/(cx+d)


How can you solve an equation with unknowns in the denominators?